1. The 2 ends of a spring balance are pulled each by a force of 10 kgwt. What will be the reading of the spring?


2. How much below the surface of the earth does the acceleration due to gravity is 10% of its value at the earth's surface ? R = 6400 km = Radius of earth.

Grade:Upto college level

1 Answers

Nehal Wani
14 years ago

Ans 1: T/=2T   Therefore, reading of spring:20N

Ans 2: gd=gs(1-d/R)

          gs/10=gs(1-d/R)

  d=9R/10=5760km